The table below shows the number of deaths in the U.S. in a year due to a variety of causes. For these questions, assume these values are not changing from year to year, and that the population of the United States is 312 million people.

Passenger car occupant 13,100
Motorcycle driver 4,500
Tornado 553
Skydiving

56

What is the probability that you will die as the result of a tornado next year? (the answer is not 553/312 million)

Why is this answer different than your answer to the previous questions?
Answer #1

Total number of deaths in the US for a particular year are 13100 + 4500 + 553 + 56 = 18209

Now define A be the event that number of deaths and B be the event define as number of deaths occur due to Tornado

we need to find P(B|A) =

= 553 / 312 million

= 18209 / 312 million

= 553 / 18209

Answer is 553/18209 = 0.03